OpenStudy (anonymous):

Show how to solve 8x^2+14x+5=0 by completing the square.

OpenStudy (anonymous):

8x^2 + 14x + 5 = 0. 8x^2 + 14x + 5 + 6 = 6. 8x^2 + 14x + 6 = 6 - 5 = 1. (4x + 3)(2x + 2) = 1 4x + 3 = 1. 4x = 1 - 3. 4x = - 2. x = - 1/2. 2x + 2 = 1. 2x = - 1. x = - 1/2. Answer.